OpenStudy (anonymous):

The length & breadth of a carboard are 11/5 and 6/5 repectively . The length and breadth of a second cardboard are 16/5 and 12/5 respectively . Both cardboards are divided into 10 equal small pieces . What is the total area of cardboard , if it is made using 5 smaller pieces from fhe 1st cardboard and 3 smaller pieces from the second cardboard ?

OpenStudy (anonymous):

I'll give medal to whom who will be answering this question😊😀☺

OpenStudy (hba):

Can you make a figure , according to the requirements of the question?

rishavraj (rishavraj):

first board area = \[\frac{ 11 }{ 5 } \times \frac{ 6 }{ 5 } = \frac{ 66 }{ 25 }\] area of one piece = \[\frac{ 66 }{ 250 }\] area of 5 such piecces = \[\frac{ 66 }{ 250 } \times 5 = \frac{ 66 }{ 50 } = \frac{ 33 }{ 25 }\]

rishavraj (rishavraj):

same steps for second board...bt u need to multiply with 3 instead of 5... :))

OpenStudy (anonymous):

Ohkkk ty so much

rishavraj (rishavraj):

u r welcome....

rishavraj (rishavraj):

area of second board = \[\frac{ 16 }{ 5 } \times \frac{ 12 }{ 5 } = \frac{ 192 }{ 25 }\] area of one small piece = \[\frac{ 192 }{ 250 }\] so area of 3 pieces = \[\frac{ 192 }{ 250 } \times 3 = \frac{ 576 }{ 250 } = \frac{ 288 }{ 125 }\]

rishavraj (rishavraj):

now area of new board = \[\frac{ 33 }{ 25 } + \frac{ 288 }{ 125 } = \frac{ 453 }{ 125 }\]

OpenStudy (anonymous):

Thank you